# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v3

@@ A.3 Does the language model use the model class we specify?
A.4 Energy distance and statistic
The energy statistic [33], which is based on the energy distance metric [50], is a method to measure the similarity of two probability distributions. Importantly, it allows us to test whether two $N$-dimensional probability distributions are equal.
@@ A.4 Energy distance and statistic
A.5 Further experimental details
In all experiments, we use OpenAI’s GPT-3.5-turbo (specifically GPT-3.5-turbo-0125) with a temperature of $0.1$ and all other settings in the OpenAI API default  [^2] https://platform.openai.com/docs/api-reference/chat.
A.5.1 Dataset details
@@
  • A private dataset collected as part of a study on Dementia care  [^3] https://www.imperial.ac.uk/uk-dri-care-research-technology/: This dataset contains $10$ features engineered from measurements of in-home activity and physiology. In this dataset, the labels corresponded to days with a clinical labelling of positive or negative urinary tract infection (UTI) which were calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of the data containing $78$ people and $371$ labelled days of UTI, with $270$ negatives and $101$ positives. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Heart Disease dataset [36, 37]: This dataset contains $10$ physiological features and targets with a positive or negative diagnosis of heart disease for $299$ people. Within this dataset, we have $161$ examples of people without heart disease and $138$ examples of people with heart disease, totalling $299$ data points. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data, except where features are categorical. In the categorical case, we perform no scaling.
  • The Hypothyroid dataset [39, 40]: This dataset, processed from the original, consists of $150$ examples of patients’ thyroid-related blood levels and whether they have a hypothyroidism diagnosis or not. This dataset contains $4$ features, with $75$ examples of a patient with hypothyroidism and $75$ examples of a patient without hypothyroidism. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Diabetes dataset [38]: After preprocessing, this dataset contains $442$ people with Diabetes and some physiological measurements. Along with some demographic information, the task is to predict a quantitative measure of their diabetes disease progression one year after this baseline. This dataset consisted of $10$ features and the targets were regression values with a maximum value of $10.0$, minimum value of $0.0$, and a mean value of $3.96$.
  • The Breast Cancer dataset [41]: This dataset is made up of $569$ data points with $30$ features computed from an image of a breast mass that are labelled as either malignant or benign. In total, this dataset contains $357$ examples of benign tumours and $212$ examples of malignant tumours. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
@@
A.5.2 Evaluating prior elicitation
To evaluate the usefulness of the priors elicited from the language model (presented in Section 2.2), we started by comparing their posterior performance to an uninformative prior. To do this, we chose the predictive tasks described in Section 2.5: $y=2x_{1}-x_{2}+x_{3}$, Urinary Tract Infection (UTI), Heart Disease, Hypothyroidism, Diabetes, and Breast Cancer and calculated the posterior accuracy or mean squared error on a test set after observing varied numbers of examples. For each of the datasets, we selected $50\%$ of data points to form the test set and used the remaining data points to form a training set. This is done $10$ times to produce $10$ folds of random training and testing splits.
@@ A.5.2 Evaluating prior elicitation
We then use the No-U-Turn sampler [35] to sample from the posterior distribution over the linear model parameters, with $5$ chains and $5000$ samples per chain. In total, for each dataset, for each training and testing split, we have $25000$ posterior samples over the posterior.
To produce the results in Figures 2(a) and 9, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. To produce Figures 2(b) and 14, we also calculate the mean accuracy and the mean squared error for each of the test splits before plotting their mean and standard deviation.
@@
A.5.3 Extracting the in-context model’s hidden prior and posterior
To further study in-context learning, we used maximum likelihood sampling to extract approximate samples of the internal model’s prior and posterior using the method described in Section 2.3.2. For each of the tasks studied, we randomly generate $25$ feature values from $X\sim\mathcal{U}[-5,5]$ $5$ times and ask the language model to predict their regression label or positive classification probability based on the task description given and with the specified model class (in our case a linear or logistic regressor). This is done for each of the $100$ task descriptions that we have for the dataset. We then fit a linear regressor to the predicted regression labels or the classification logits  [^4] Calculated using the inverse logistic function: $\text{logits}=\ln p/(1-p)$ of the predicted probabilities, using maximum likelihood estimation. The parameters of this linear model are then taken as a single sample of the approximated in-context predictive model.
@@ A.5.3 Extracting the in-context model’s hidden prior and posterior
To do this, we start by fitting a Gaussian kernel density estimator (with a bandwidth factor of $0.25$) to our in-context model prior parameter samples. We then use the No-U-Turn sampler [35] with $1000$ adaption steps and the same demonstration examples given during in-context learning to sample posterior parameters. Here, we use $100$ chains of $10000$ samples to build an approximate posterior distribution.
@@
A.5.4 Calculating Bayes factor
@@ A.5.5 Given a new dataset
A.6 Further prior elicitation results
In Figure 9 we present the results of the same experiment as discussed in Section 3.1, applied to three other datasets discussed in Section 2.5.
For the synthetic task $y=2x_{1}-x_{2}+x_{3}$, the priors elicited by LLM allow for a posterior predictive distribution with a markedly reduced mean squared error over an uninformative prior and often produce less variation in the result. Due to the simple nature of the task however, after only a few training examples, both priors achieve similar posterior results.
In the case of Hypothyroid prediction, the LLM-elicited prior produces similarly accurate posterior parameter samples than the uninformative prior for smaller numbers of observations. However, as the number of observations increases, the LLM-elicited prior begins to produce more accurate posterior distributions.
For the Heart Disease dataset, we find that the prior elicited from the language model performed similarly to the uninformative prior. This is because the language model provided mean = 0 and std = 1 for a large number of priors elicited, suggesting that the language model was unsure about the prior distribution on the parameters for a model predicting heart disease. The parameter distributions elicited can be seen in Figure 12, which contextualises these results.
A.6.1 Priors elicited from the synthetic data
Figure 10 shows a histogram of parameter values from the GPT-3.5-turbo elicited prior for the task descriptions of $y=2x_{1}-x_{2}+x_{3}$. Here, the language model correctly provides prior distributions that are centred around the correct parameter values and offer a small standard deviation.
A.6.2 UTI accuracy with the label collection context
Figure 11: Posterior accuracy of the elicited priors on the UTI dataset considering label collection time. The average mean posterior accuracy and 95% confidence interval on the UTI test set after observing the given number of data points against the time it took for the data points to be collected in the study.
In Figure 11 we plot the time it took to collect the given number of labels tested in the experiments in Section 3.1 against the accuracy of a model using an uninformative prior and one elicited from a language model. This result illustrates the usefulness of prior elicitation, since we are able to achieve comparable (and ultimately better) accuracy with fewer data points, and therefore less data collection time.
Since UTI label collection is costly and slow, our UTI predictive model could have achieved greater accuracy significantly earlier in the study. This is paramount in a clinical environment, where an improved UTI detection model could have resulted in fewer unplanned hospital admissions.
A.7 Is the language model being truthful when we elicit priors?
Figure 12: Distribution of the prior parameters under prior elicitation and those extracted from the internal model for the different predictive tasks. Here, we compare the prior distributions of the different predictive tasks, as defined by the internal model used in-context, and that which we elicit from the language model. Here, we draw the histogram separately in each dimension, but calculate the energy distance across all features combined.
In Section 3.4, we study how truthful the language model is when we elicit a prior distribution, compared to the prior it uses in its internal predictive model when performing in-context learning.
A.8 Empirically, is the language model approximating Bayesian inference?
In this section, we present the differences between the posterior distribution extracted through maximum likelihood sampling of the LLM’s in-context internal model (Section 3.3) and one that is sampled through MC methods on the extracted in-context prior distribution.
Table 2: Difference between extracted posterior and MC on extracted prior. The energy statistic between the language model’s internal posterior extracted through MLE sampling and the posterior samples from applying Monte Carlo techniques to the language model’s internal prior. Values are $\in[0,1]$ with values closer to zero, indicating that the two distributions are more similar. Here, we show the mean and standard deviation of the energy statistic for the $5$ different training sets tested as observed data.
Heart Disease Hypothyroid Diabetes Breast Cancer
0.04 ± 0.01 0.06 ± 0.01 0.02 ± 0.00 0.01 ± 0.00
Figure 13: The different posterior samples Here, we compare the distribution of the posterior extracted from the language model’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined.
A.9 Dataset memorisation
To better contextualise the use of in-context learning in the tested datasets, we will use the method proposed in Bordt et al. [42] to quantify the extent to which the language model already knows information about the datasets.
@@ A.9 Dataset memorisation
We then use the Levenshtein distance [51, 52], a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00, suggesting that the strings are equal, and a maximum value of the longest length of the two strings.
@@
Table 3: Header completion test results. Normalised Levenshtein distance between the correct header completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
@@
Table 4: Row completion test results. Normalised Levenshtein distance between the correct row completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
@@
A.9.1 How does this affect our results?
@@
A.10 Task descriptions with more information
@@
Figure 14: Mean squared error using the elicited priors for the task: $y=2x_{1}-x_{2}+x_{3}$ where we have used increasingly more detailed task descriptions. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$. Within each task description we provide the language model with increasingly more information about the task, as to test how the task description effects the priors elicited. Complete descriptions of the task descriptions given here are provided in Section A.10. The lines depict the average mean posterior mean squared error (MSE) over the 5 dataset splits, whilst the error bars represent the 95% confidence interval.
@@
To understand how the detail of the task description affects the performance of the priors elicited from the language model, we studied the synthetic task, $y=2x_{1}-x_{2}+x_{3}$, and described the relationship between features and targets in varied detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 14, we included the following text in the task description that was provided to the language model:
@@
However, what is unexpected is that for this task, using an uninformative prior ($\theta\sim\mathcal{N}(\theta|0,1)$) provides better posterior performance than specifying to the language model that the relationship between the target and features is linear (Figure 14). This is explained by inspecting the prior distribution elicited when given this minimal information, which shows that the language model responded with a similar mean and standard deviation for a Gaussian for each of the $100$ variations of the prompts:
@@
This demonstrates the sometimes unexpected behaviour of language models, which here produced Gaussians for the prior distribution with small standard deviations (reflecting a high confidence in the distribution), and means that imply the language model knew the underlying equation defining the relationship between the target and features. We might have expected that after only being told the relationship was linear, the language model would produce Gaussians with 00 mean and standard deviation $1$, reflecting the uncertainty in the exact relationship between each feature and the target.
@@
Figure 15: Parameter distribution as further feature information is included. These histograms show the distribution of parameter values for the features as more expert information is provided. In each row, we provide information in the task description of an additional feature, with the green arrow indicating the effect this has on the parameter distribution. In particular, in the first row we provide information about the relationship between the target and the first feature, in the second row we provide information about the second feature, and the in the third row we provide information about the third feature.
An expert interacting with the language model might be able to provide further information about a task that improves the elicited priors. In Figure 15 we show the parameter distributions elicited from the language model for the different descriptions we provide. This figure shows that, as we provide additional information about each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. This demonstrates the potential of using an expert to construct more detailed task descriptions, which the language model can translate into parameter distributions. This experiment provides a controlled setting that supplements the results in Section 3.2.
A.11 How many ways should we describe a task?
Figure 16: Measuring how similar the prior distributions are with fewer task descriptions. We reduce the number of task descriptions and calculate the energy statistic between $10000$ samples from the prior elicited from the language model using the reduced set of descriptions and the full $100$ task descriptions. We repeat each experiment $10$ times and show the mean and standard deviation.
In this section, we look at how many of the task descriptions are required to be able to get a similar prior distribution to that elicited within our experiments, using $100$ task descriptions.

### v4

@@ A.4 Does the language model use the model class we specify?
A.5 Energy distance and statistic
The energy statistic (Székely & Rizzo, 2013), which is based on the energy distance metric (Székely & Rizzo, 2012), is a method to measure the similarity of two probability distributions. Importantly, it allows us to test whether two $N$-dimensional probability distributions are equal.
@@ A.5 Energy distance and statistic
A.6 The Bayes factor
Given two models $\alpha_{0}$ and $\alpha_{1}$ with parameterisations $\theta_{0}$ and $\theta_{1}$, the Bayes factor calculated over a dataset $\mathcal{D}$ is:
[…]
In our case, $\alpha_{0}$ refers to prior elicitation, with the prior determining the parameterisation $\theta$ of a predictive model, and $\alpha_{1}$ referring to in-context learning. We can approximate the Bayes factor using the means from samples of the likelihood of $D$.
A.7 Further experimental details
In all experiments, we use OpenAI’s GPT-3.5-turbo (specifically GPT-3.5-turbo-0125), GPT-4-turbo (GPT-4-turbo-2024-04-09) [^2] https://platform.openai.com/docs/api-reference/chat, and DeepSeek-R1-32B-Q4333https://ollama.com/library/deepseek-r1:32b with a temperature of $0.1$ and all other settings as default.
A.7.1 Dataset details
@@
  • A private dataset collected as part of our study on Dementia care (Capstick et al., 2024) [^4] https://www.imperial.ac.uk/uk-dri-care-research-technology/: This dataset contains $10$ features engineered from measurements of in-home activity and physiology. The raw features were measured using passive infrared (PIR) sensors located around the homes of people living with dementia, and a sleep mat positioned in the participants bed. The targets correspond to days with a clinical labelling of positive or negative urinary tract infection (UTI) which were calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of the data containing $78$ people and $371$ labelled days of UTI, with $270$ negatives and $101$ positives. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Heart Disease dataset (Detrano et al., 1989; Andras Janosi, 1989): This dataset contains $10$ physiological features and targets with a positive or negative diagnosis of heart disease for $299$ people. Within this dataset, we have $161$ examples of people without heart disease and $138$ examples of people with heart disease, totalling $299$ data points. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data, except where features are categorical. In the categorical case, we perform no scaling.
  • The Hypothyroid dataset (Quinlan, 1986; Turney, 1994): This dataset, processed from the original, consists of $150$ examples of patients’ thyroid-related blood levels and whether they have a hypothyroidism diagnosis or not. This dataset contains $4$ features, with $75$ examples of a patient with hypothyroidism and $75$ examples of a patient without hypothyroidism. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Diabetes dataset (Efron et al., 2004): After preprocessing, this dataset contains $442$ people with Diabetes and some physiological measurements. Along with some demographic information, the task is to predict a quantitative measure of their diabetes disease progression one year after this baseline. This dataset consisted of $10$ features and the targets were regression values with a maximum value of $10.0$, minimum value of $0.0$, and a mean value of $3.96$.
  • The Breast Cancer dataset (Wolberg et al., 1995): This dataset is made up of $569$ data points with $30$ features computed from an image of a breast mass that are labelled as either malignant or benign. In total, this dataset contains $357$ examples of benign tumours and $212$ examples of malignant tumours. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
@@
A.7.2 Language model details
Within the main experiments, we evaluate three language models:
  • GPT-3.5-turbo [^5] https://platform.openai.com/docs/models is one of OpenAI’s least expensive models and is described as a model for simple tasks.
  • GPT-4-turbo [^6] https://platform.openai.com/docs/models is one of OpenAI’s more sophisticated models.
  • DeepSeek-R1-32B-Q4777https://ollama.com/library/deepseek-r1:32b, a 32 billion parameter model from DeepSeek based on Qwen 32B and distilled from DeepSeek-R1.
A.7.3 Evaluating LLM-elicited priors
To evaluate the usefulness of the AutoElicit priors (presented in Section 4.1), we started by comparing their posterior performance to an uninformative prior. To do this, we chose the predictive tasks described in Section 5.1: $y=2x_{1}-x_{2}+x_{3}$, Urinary Tract Infection (UTI), Heart Disease, Hypothyroidism, Diabetes, and Breast Cancer and calculated the posterior accuracy or mean squared error on a test set after observing varied numbers of examples. For each of the datasets, we selected $50\%$ of data points to form the test set and used the remaining data points to form a training set. This is done $10$ times to produce $10$ folds of random training and testing splits.
@@ A.7.3 Evaluating LLM-elicited priors
We then use the No-U-Turn sampler (Hoffman et al., 2014) to sample from the posterior distribution over the linear model parameters, with $5$ chains and $5000$ samples per chain. In total, for each dataset, for each training and testing split, we have $25000$ posterior samples over the posterior.
To produce the results in Figure 9, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. To produce Figures 2 and 12, we also calculate the mean accuracy and the mean squared error for each of the test splits before plotting the mean and the $95\%$ confidence interval.
@@
A.7.4 Extracting the in-context model’s hidden prior and posterior
To further study in-context learning, we used maximum likelihood sampling to extract approximate samples of the internal model’s prior and posterior using the method described in Section 4.2. For each of the tasks studied, we randomly generate $25$ feature values from $X\sim\mathcal{U}[-5,5]$ $5$ times and ask the language model to predict their regression label or positive classification probability based on the task description given and with the specified model class (in our case a linear or logistic regressor). This is done for each of the $100$ task descriptions that we have for the dataset. We then fit a linear regressor to the predicted regression labels or the classification logits [^8] Calculated using the inverse logistic function: $\text{logits}=\ln p/(1-p)$ of the predicted probabilities, using maximum likelihood estimation. The parameters of this linear model are then taken as a single sample of the approximated in-context predictive model.
@@ A.7.4 Extracting the in-context model’s hidden prior and posterior
To do this, we start by fitting a Gaussian kernel density estimator (with a bandwidth factor of $0.25$) to our in-context model prior parameter samples. We then use the No-U-Turn sampler (Hoffman et al., 2014) with $1000$ adaption steps and the same demonstration examples given during in-context learning to sample posterior parameters. Here, we use $100$ chains of $10000$ samples to build an approximate posterior distribution.
@@
A.7.5 Calculating Bayes factor
@@ A.7.6 Given a new dataset
A.8 Dataset memorisation
To better contextualise the use of in-context learning in the tested datasets, we will use the method proposed in Bordt et al. (2024) to quantify the extent to which GPT-3.5-turbo already knows information about the datasets.
@@ A.8 Dataset memorisation
We then use the Levenshtein distance (Levenshtein, 1965; 1966), a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00, suggesting that the strings are equal, and a maximum value of the longest length of the two strings.
@@
Table 2: GPT-3.5-turbo header completion test results. Normalised Levenshtein distance between the correct header completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
@@
Table 3: GPT-3.5-turbo row completion test results. Normalised Levenshtein distance between the correct row completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
@@
A.8.1 How does this affect our results?
@@
A.9 Additional AutoElicit results
In Figure 9 we present the accuracy distributions from the experiment in Figure 2, from Section 5.2.
For the synthetic task $y=2x_{1}-x_{2}+x_{3}$, the priors elicited from the LLMs allow for a posterior predictive distribution with a markedly reduced mean squared error (MSE) over an uninformative prior whilst producing less variation in the result. Here, GPT-3.5-turbo initially produces the lowest MSE. However, due to the simple nature of the task, as more data becomes available, all priors achieve similar results.
When predicting UTI, all of the AutoElicit priors provide improved accuracy over the uninformative prior for all training sizes tested. This should be of particular interest, given that the UTI dataset is private and so will not have appeared in any form in the training data for the LLM.
For the Heart Disease dataset, we find that the prior elicited from the language model performed similarly to the uninformative prior. This is because the language model provided mean = 0 and std = 1 for a large number of priors elicited, suggesting that the LLM was unsure of its prior knowledge. The parameter distributions elicited can be seen in Figure 15, which contextualises these results. However, DeepSeek-R1-32B-Q4 and GPT-4-turbo provided prior distributions that were considerably more accurate.
For Hypothyroid prediction, DeepSeek-R1-32B-Q4 provides the best accuracy by a large margin. Of the OpenAI models, smaller training sizes favour GPT-4-turbo, whilst larger training sizes favour, GPT-3.5-turbo. This is because the GPT-4-turbo-elicited prior was too strict to allow for updated posterior parameters, but was more informative than an uninformative prior for smaller training sizes.
For the Diabetes task, both OpenAI LLMs provide considerably improved MSE over the uninformative prior for training sizes up to $50$ observed data points, whilst DeepSeek-R1-32B-Q4 provides worse MSE.
In the Breast Cancer task, the accuracy improvement from the elicited DeepSeek-R1-32B-Q4 and GPT-4-turbo priors is maintained across all training sizes, whilst GPT-3.5-turbo provides an improvement up to $40$ observed data points. This can be seen in these distributions, which show a long tail in the accuracy distribution for the uninformative prior for most training sizes.
A.9.1 Priors elicited from the synthetic data
Figure 10 shows the histogram of parameter values from the GPT-3.5-turbo-elicited prior for the task descriptions of $y=2x_{1}-x_{2}+x_{3}$. Here, the language model correctly provides prior distributions that are centred around the correct parameter values and offer a small standard deviation.
A.9.2 UTI accuracy with the label collection context
Figure 11: Posterior accuracy of the elicited priors on the UTI dataset considering label collection time. The average mean posterior accuracy and 95% confidence interval on the UTI test set after observing the given number of data points against the time it took for the data points to be collected in the study. The horizontal axis shows the time in days to achieve the given performance and the number of labels collected in that time.
In Figure 11 we plot the time it took to collect the given number of labels tested in the experiments in Section 5.2 against the accuracy of a model using an uninformative prior and one elicited from a language model. This result illustrates the usefulness of prior elicitation, since we are able to achieve a greater accuracy with fewer data points, and therefore less data collection time.
Since UTI label collection is costly and slow, our predictive model could have achieved greater accuracy significantly earlier in the study. This is paramount in a clinical environment, where an improved UTI detection model could have resulted in fewer unplanned hospital admissions.
A.10 Cost of AutoElicit
The priors elicited from GPT-3.5-turbo and GPT-4-turbo incur a cost, as we use OpenAI’s API. To elicit all of the priors used in the experiments in Section 5.2, we were charged \$$\$0.41$ (on average \$$\$0.068$ per dataset) for GPT-3.5-turbo and \$$\$7.56$ (on average \$$\$1.26$ per dataset) for GPT-4-turbo. As a rule, the more features a dataset contains or the longer the task description, the more expensive priors are to elicit since the number of input and output tokens is greater. However, we did not make use of the batch API and instead made each API call sequentially. All experiments in this work could be implemented using the batch API since no new API call depends on a previous API result, which may significantly reduce costs [^9] https://openai.com/api/pricing/.
Eliciting the DeepSeek-R1-32B-Q4-based prior distributions incurred no API cost, because the model is open-source and we were able to run it locally. However, to do this, we used an Nvidia A100, which will likely not be available in a low-resource setting. The API costs for the non-quantised model will apply in these cases [^10] https://api-docs.deepseek.com/quick_start/pricing.
A.11 Task descriptions with more information
@@
Figure 12: Mean squared error using the GPT-3.5-turbo-elicited priors for the task: $y=2x_{1}-x_{2}+x_{3}$ where we have used increasingly more detailed task descriptions. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$. Within each task description we provide the language model with increasingly more information about the task, as to test how the task description effects the priors elicited. Complete descriptions of the task descriptions given here are provided in Section A.11. The lines depict the average mean posterior mean squared error (MSE) over the 5 dataset splits, whilst the error bars represent the 95% confidence interval.
@@
To understand how the detail of the task description affects the performance of the priors elicited from the language model, we studied the synthetic task $y=2x_{1}-x_{2}+x_{3}$, and describe the relationship between features and targets in increasing detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 12, we included the following text in the task description that was provided to the language model:
@@
However, it is unexpected that using an uninformative prior ($\theta\sim\mathcal{N}(\theta|0,1)$) provides better posterior performance than specifying to the language model that the relationship between the target and features is linear (Figure 12). This is explained by inspecting the prior distribution elicited when given this minimal information, which shows that the language model responded with a similar mean and standard deviation for a Gaussian for each of the $100$ variations of the prompts:
@@
This demonstrates the sometimes unexpected behaviour of language models. Here, the LLM produced Gaussians for the prior distribution with small standard deviations (reflecting a high confidence in the distribution) and means that imply the language model knew the underlying equation defining the relationship between the target and features. We might have expected that after only being told the relationship was linear, the language model would produce Gaussians with 00 mean and standard deviation $1$, reflecting the uncertainty in the exact relationship between each feature and the target.
@@
Figure 13: GPT-3.5-turbo-elicited parameter distribution as increasingly more information is included. These histograms show the distribution of parameter values for the features as more expert information is provided. In each row, we provide information in the task description of an additional feature, with the green arrow indicating the effect this has on the parameter distribution. In particular, in the first row we provide information about the relationship between the target and the first feature, in the second row we provide information about the second feature, and the in the third row we provide information about the third feature.
An expert interacting with the language model might be able to provide further information about a task that improves the elicited priors. In Figure 13 we show the parameter distributions elicited from the language model for the different descriptions we provide. As we provide additional information on each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. This demonstrates the potential of using an expert to construct more detailed task descriptions, which the language model can translate into parameter distributions. This experiment provides a controlled setting that supplements the results in Section 5.3.
A.12 How many ways should we describe a task?
Figure 14: Measuring how similar the GPT-3.5-turbo-elicited prior distributions are with fewer task descriptions. We reduce the number of task descriptions and calculate the energy statistic between $10000$ samples from the prior elicited from the language model using the reduced set of descriptions and the full $100$ task descriptions. We repeat each experiment $10$ times and show the mean and standard deviation.
In this section, we examine how many task descriptions are required to be able to get a similar prior distribution to that elicited within our experiments, using $100$ task descriptions.